OpenStudy (anonymous):

The function f(x) varies directly with x and f(x) = 250 when x = 50. What is f(x) when x = 5?

OpenStudy (anonymous):

Since there is a direct relation, so f(x)=k*x. Now, putting the values in the relation 250=k*50 i.e. k=5. Hence for x=5, f(x)=5*5=25